Calculate the quantity of heat absorbed by 3 kg of copper when heated from room temperature to 70oC. (ccu= 0.092 cal/groC)




1
Expert's answer
2021-02-07T19:10:38-0500
"Q=mc\Delta T,""Q=3000\ g\cdot0.092\ \dfrac{cal}{g\cdot \!^{\circ}C}\cdot(70\ ^{\circ}C-20\ ^{\circ}C)=13800\ cal=13.8\ kcal."